### v1

# WAFT: Warping-Alone Field Transforms for Optical Flow

Abstract
We introduce Warping-Alone Field Transforms (WAFT), a simple and effective method for optical flow. WAFT is similar to RAFT but replaces cost volume with high-resolution warping, achieving better accuracy with lower memory cost. This design challenges the conventional wisdom that constructing cost volumes is necessary for strong performance. WAFT is a simple and flexible meta-architecture with minimal inductive biases and reliance on custom designs. Compared with existing methods, WAFT ranks 1st on Spring and KITTI benchmarks, achieves the best zero-shot generalization on KITTI, while being up to 4.1×4.1$\times$4.1 × faster than methods with similar performance. Code and model weights are available at https://github.com/princeton-vl/WAFT.
1 Introduction
Optical flow is a fundamental level vision task that estimates per-pixel 2D motion between video frames. It has many downstream applications, including 3D reconstruction and synthesis [29, 68], action recognition [49, 38, 65], frame interpolation [59, 25, 14], and autonomous driving [9, 33, 19].
Cost volumes [46, 16] with iterative updates [50, 52] has become a standard design in most state-of-the-art methods [46, 8, 58, 50, 13, 52, 34], especially when both accuracy and efficiency are taken into account. Previous work [46, 50] regards cost volume as a more effective representation than image features, as it explicitly models the visual similarity between pixels.
However, constructing cost volumes is expensive in both time and memory [64, 55]. The cost increases quadratically with the radius of the neighborhood. As a result, cost volumes are often constructed from low resolution features, limiting the ability of the model to handle high-resolution input images.
In this paper, we challenge the conventional wisdom that cost volume is necessary for strong performance with high efficiency, and introduce Warping-Alone Field Transforms (WAFT), a simplified design that replaces cost volumes with warping and achieves state-of-the-art accuracy across various benchmarks with high efficiency.
For each pixel in frame 1, instead of computing its similarities against many pixels in frame 2, warping simply fetches the feature vector of the corresponding pixel given by the current flow estimate; this enables memory-efficient high-resolution processing and leads to better accuracy.
The design of WAFT is simple, with flow-specific designs kept to the minimum. WAFT consists of an input encoder that extracts features from individual input frames and a recurrent update module that iteratively updates flow. Compared to other RAFT-like architectures, WAFT is much simplified because it does not use cost volumes and has removed the context encoder that provides extra features for the update module. WAFT is designed to function as a meta-architecture for optical flow in the sense that the individual components including input encoder and the update unit do not require custom designs and can use existing off-the-shelf (pretrained) architectures. In our experiments, we evaluate different choices (such as ResNet [11] and DPT [39]) that have different accuracy-efficiency trade-offs.
WAFT achieves state-of-the-art performance across various benchmarks with high efficiency and a simple design. On KITTI [9], WAFT achieves the best Fl on non-occluded pixels and the second best on all pixels. On Spring [31], WAFT ranks first across all metrics. It also achieves the best zero-shot cross-dataset generalization on KITTI. We achieve this with standard network architectures for the sub-modules  [7, 11, 39], removing custom designs typically needed in prior work, while being up to 4.1$\times$ faster than existing methods.
Our main contributions are two fold: (1) we challenge the conventional wisdom that cost volume is a key component for achieving state-of-the-art accuracy and efficiency for optical flow; (2) we introduce WAFT, a warping-based meta-architecture that is simpler and achieves state-of-the-art accuracy with high efficiency.
Figure 1: The meta-architecture of WAFT consists of an input encoder and a recurrent update module. We first extract image features from the input encoder, and then use these features to iteratively update the flow estimate for $T$ steps. At each step, we perform feature indexing through a lightweight backward warping on the feature of frame 2, removing the dependency on expensive cost volume used by previous work [46, 50].
2 Related Work
Estimating Optical Flow
Traditional methods treated optical flow as a global optimization problem that maximizes visual similarity between corresponding pixels [12, 61, 4, 2]. These methods apply coarse-to-fine warping [2, 1, 32], a strategy theoretically justified by Brox et al. [2], to solve this optimization.
Today, this field is dominated by deep learning methods [16, 8, 46, 62, 15, 50, 45, 48, 5, 13, 44, 53, 54, 56, 57, 24, 43, 17, 18, 27, 66, 63, 28, 21, 26, 67, 34], which can be categorized into two paradigms: direct or iterative. Direct methods [8, 53, 54, 43, 24, 56] treat flow estimation as a standard dense prediction task (e.g. monocular depth estimation) and directly regress the dense flow field from large-scale pre-trained models. Iterative methods [50, 52, 26, 34, 67, 46, 13] align more closely with traditional warping-based approaches, refining the flow predictions progressively. Most state-of-the-art methods [50, 52, 13, 44, 26, 34] follow the iterative paradigm due to its significantly higher efficiency than the direct ones.
Cost volumes [46, 8] have been regarded as a standard design in iterative methods [50, 52, 34, 13, 44, 57, 26]. Prior work [50, 46] empirically shows the effectiveness of cost volumes in handling large displacements. Many iterative methods [50, 52, 34] adopt partial cost volumes [46] to avoid the quadratic computational complexity of full 4D cost volumes; they restrict the search range of each pixel in frame 1 to the neighborhood of its corresponding pixel in frame 2. However, they still suffer from the high memory consumption inherent in cost volumes [55, 64].
WAFT is a warping-based iterative method. We achieve state-of-the-art performance across various benchmarks without constructing cost volumes, challenging the conventional wisdom established by previous work [46, 50, 13]. Warping no longer suffers from high memory consumption inherent in cost volumes, which enables high-resolution indexing and therefore improves accuracy.
Vision Transformers
Vision transformers [7] have achieved significant progress across a wide range of visual tasks [60, 22, 37, 10, 42]. In the context of optical flow, most direct methods [53, 54, 43] regress flow from a large-scale pre-trained vision transformer with a lightweight flow head. Many iterative methods [13, 44, 26, 67] design task-specific transformer blocks to process cost volumes.
WAFT adopts similar designs to DPT [39] in its recurrent update module, which implicitly handles large displacements in optical flow through the transformer architecture. We empirically show that this is crucial to make warping work. WAFT can also benefit from large-scale pre-trained transformers like existing methods [43, 53, 54, 67], with minimal additional flow-specific designs.
3 Background
In this section, we first review current cost-volume-based iterative methods and discuss the drawbacks of cost volumes. Then we introduce warping and compare it to cost volumes.
Figure 2: For each pixel, the full cost volume calculates its visual similarity to all pixels in the other frame through correlation. The partial cost volume restricts the search range to the neighborhood of the corresponding pixel, marked by a red box. Compared with them, warping only uses the information from the corresponding pixel, offering better time and memory efficiency. This efficiency enables high-resolution processing, which leads to improved accuracy.
3.1 Iterative Methods with Cost Volume
Given two adjacent RGB frames, optical flow predicts pixel-wise 2D motion between adjacent frames. Current iterative methods [50, 13, 52] consist of two parts: (1) input encoders that extract dense image features at low resolution, and (2) a recurrent update module that iteratively refines the flow estimate.
Denoting the two frames as $I_{1},I_{2}\in\mathbb{R}^{H\times W\times 3}$, the input encoder $F$ maps $I_{1}$, $I_{2}$ to low-resolution dense features $F(I_{1}),F(I_{2})\in\mathbb{R}^{h\times w\times d}$, respectively. A cost volume $V\in\mathbb{R}^{h\times w\times h\times w}$ is built on these features, which explicitly models the correlation between pixels ($p\in I_{1}$, $p^{\prime}\in I_{2}$):
$V_{p,p^{\prime}}=F(I_{1})_{p}\cdot F(I_{2})_{p^{\prime}}$
where $\cdot$ represents the dot product of two vectors. At each step, the recurrent update module indexes into the cost volume using the current flow estimate and predicts the residual flow update. Several methods [13, 44] directly process $V$ for better performance, but tend to be more costly.
Partial cost volume [46] is introduced to reduce the cost by avoiding the construction of a full 4D cost volume. Given the current flow estimate $f_{cur}\in\mathbb{R}^{h\times w\times 2}$ and a pre-defined look-up radius $r$, partial cost volume $V_{par}:\mathbb{R}^{h\times w\times 2}\to\mathbb{R}^{h\times w\times(r^{2})}$ implements an on-the-fly partial construction by restricting the indexing range of a pixel $p\in I_{1}$ to the neighborhood of its corresponding pixel $p+(f_{cur})_{p}\in I_{2}$, formulated as:
$V_{par}(f_{cur};r)_{p}=\text{concat}(\{V_{p,p^{\prime}}|\forall p^{\prime}\in I_{2},\text{s.t.}\|p+(f_{cur})_{p}-p^{\prime}\|_{\infty}\leq r\|)$
where the operator “concat” concatenates all values inside the set into a vector. In practice, partial cost volumes are usually constructed at multiple scales [50, 52] to improve the prediction of large displacements. Current methods also introduce context encoder [46, 50, 13, 52] to enhance the effectiveness of iterative refinement.
3.2 Drawbacks of Cost Volumes
Figure 3: Visualizations of different methods [52, 34] on Spring [31]. WAFT, benefiting from high-resolution indexing, obtains sharper boundaries and lower errors than low-resolution approaches.
Method Training Memory Cost (GiB)
1/8 Reso. 1/4 Reso. 1/2 Reso.
SEA-RAFT [52] 14.1 25.8 OOM
WAFT (Ours) - - 9.5
Table 1: We profile the training memory cost with batch size 1 on an RTX A6000. Our warping method significantly reduces the cost.
High Memory Cost
The main drawback of cost volume is its high memory consumption. Full or partial cost volume at high resolution are very expensive and often infeasible [64, 55]. Therefore, most iterative methods build the cost volume and index into it at 1/8 resolution. To further demonstrate the problem, we implement several variants of SEA-RAFT [52] that build partial cost volumes at different resolutions. We set the base channel dimension as 32, 64, and 128 for the 1/2, 1/4, and 1/8 resolution variants, respectively, to make their computational cost similar (around 350GMACs). As shown in Table 1, the partial cost volume with look-up radius $r=4$ in SEA-RAFT runs out of memory at 1/2 resolution.
Error from Low Resolution Indexing
Since cost volumes are restricted to low resolution, the predicted flow field must be downsampled for cost volume look-up, which inevitably introduces errors. As illustrated in Figure 3 using an example from the Spring benchmark [31], existing methods struggle to produce clear boundaries, particularly noticeable in the top-right corner. In contrast, our method benefits from high-resolution look-up of feature vectors and obtains sharper boundary predictions in these challenging regions. We also show the quantitative performance gain from high-resolution indexing in Table 5.
3.3 Warping vs. Cost Volume
Warping was widely used in both classical and early deep learning approaches [32, 2, 16, 40]. However, most recent iterative methods [52, 50, 13] have either replaced warping with cost volumes or used both in combination, since cost volumes have been shown to remarkably improve the performance [46]. In this section, we analyze the similarities and differences between warping and cost volume (see Figure 2), and argue that with appropriate designs, warping-based methods can achieve performance on par with cost-volume-based methods.
The overlap of warping and cost volumes lies in their use of the current flow prediction to index into feature maps, which is closely related to optimization [2]. In cost-volume-based iterative methods, the current flow estimate $f_{cur}$ is used to define partial cost volume $V_{par}$ (see Section 3.1). It is also used to define the warped feature map $\texttt{Warp}(f_{\text{cur}})\in\mathbb{R}^{h\times w\times d}$, where the feature vector of pixel $p\in I_{1}$ is indexed from the feature map of frame 2, formulated as:
$\texttt{Warp}(f_{\text{cur}})_{p}=F(I_{2})_{p+(f_{\text{cur}})_{p}}$
Compared to cost volumes, for each pixel in frame 1, warping does not calculate its visual similarity to multiple pixels in frame 2, making it no longer able to explicitly model large displacements. However, we can implicitly handle this long-range dependence through the attention mechanism in vision transformers [7, 39], which, as we will demonstrate, is crucial to make warping work well (see Section 5). The high memory efficiency of warping also enables high-resolution indexing, leading to improved accuracy.
4 Method
Figure 4: Visualizations on Spring [31], KITTI [9], and Sintel [3].
In this section, we describe WAFT, our warping-based iterative method, shown in Figure 1. Its design can be understood as a simple meta-architecture that integrates an input encoder and a recurrent update module. We will also discuss the advantages of our design, especially on its strong performance and the simplifications over past designs.
Input Encoder
The input encoder $F$ takes a single image $I$ as input and outputs a feature map. The input encoder can be based on current large-scale pre-trained models. We use frozen features from DepthAnythingV2 [60], and further refine them with a ResNet [11]. Specifically, we concatenate the feature from DepthAnythingV2 [60] with the original image and feed it into the ResNet [11], followed by a deconvolution head which upsamples the low-resolution ResNet feature to half of the original image resolution: i.e. $(h,w)=(\frac{H}{2},\frac{W}{2}),F(I)\in\mathbb{R}^{h\times w\times d}$.
Recurrent Update Module
Similar to existing iterative methods [50, 52, 13], our recurrent update module $R$ iteratively predicts the residual flow updates. At step $t$, we concatenate $F(I_{1})$ (feature of frame 1), $\texttt{Warp}(f_{\text{cur}})$ (warped feature of frame 2, see Section 3.3), and the current hidden state $\text{Hidden}_{t}\in\mathbb{R}^{h\times w\times d}$ as input. We use a slightly modified DPT [39] as the architecture of the module.
Prediction Head & Loss
We adopt the Mixture-of-Laplace (MoL) loss used in SEA-RAFT [52]. At step $t$, the hidden state $\text{Hidden}_{t}$ is used to predict the MoL parameters $M\in\mathbb{R}^{h\times w\times 6}$. They are upsampled to the original image resolution through convex upsampling [50, 52].
Simplifications over Existing Iterative Methods
We replace cost volumes, which are standard in existing iterative methods [46, 50, 13, 52], with high-resolution warping, which is more memory-efficient. In addition, we have removed the context encoder [46], another flow-specific design standard in existing iterative methods.
A direct benefit from our simplified design is that we can load pre-trained weights for standard architectures such as DPT [7, 39], which can improve generalization as our experiments will show Section 5.
The simplicity of our meta-architecture also enables more apples-to-apples comparisons between direct methods and iterative methods. Existing direct methods [53, 54, 43] share an input format similar to that of the first iteration of WAFT, making them more directly compatible. We will empirically show the effectiveness and necessity of iterative indexing within our meta-architecture in Section 5.
5 Experiments
Type Method Sintel KITTI Inference Cost
[…]
Direct GMFlow [56] 1.74 2.90 9.32 3.80 603 139
CroCoFlow [54] 1.09 2.44 3.64 2.40 57343 6422
DDVM [43] 1.75 2.48
3.26
2.24 - -
Iterative w/ Cost Volume PWC-Net+ [47] 3.45 4.60 7.72 4.91 101 24
RAFT [50] 1.61 2.86 5.10 3.07 938 141
DIP [66] 1.44 2.83 4.21 2.43 3068 499
GMFlowNet [63] 1.39 2.65 4.79 2.75 1094 244
CRAFT [45] 1.45 2.42 4.79 3.02 2274 483
FlowFormer [13] 1.20 2.12 4.68 2.69 1715 336
GMFlow+ [57] 1.03 2.37 4.49 2.40 1177 250
RPKNet [35] 1.31 2.65 4.64 2.71 137 183
CCMR+ [17] 1.07 2.10 3.86 2.07 12653 999
MatchFlow(G) [6] 1.16 2.37 4.63 2.77 1669 291
Flowformer++[44] 1.07
1.94
4.52 - 1713 374
SEA-RAFT(L) [52] 1.31 2.60 4.30 - 655 108
AnyFlow [21] 1.23 2.44 4.41 2.69 - -
FlowDiffuser [26] 1.02 2.03 4.17 2.82 2466 599
SAMFlow [67]
1.00
2.08 4.49 - 9717 1757
DPFlow [34] 1.04 1.97 3.56 2.12 414 131
Iterative w/ Warping SpyNet [40, 36] 6.64 8.36 35.07 26.71 167 25
FlowNet2 [16] 4.16 5.74 10.41 6.94 230 75
WAFT (Ours)
1.09 2.34 3.42
2.04
853 240
Table 2: We report endpoint-error (EPE) on Sintel [3] and Fl on KITTI [9]. On KITTI [9], WAFT ranks first on non-occluded pixels and second on all pixels. It also achieves comparable performance to state-of-the-art methods on Sintel (clean). We measure latency on an RTX3090 with batch size 1 and input resolution $540\times 960$.
We report results on Sintel [3], KITTI [9], and Spring [31]. Following existing work [50, 13, 52, 34], for training, we use FlythingChairs [8], FlyingThings [30], HD1K [23], Sintel [3], KITTI [9], Spring [31], and TartanAir [51]. We use the widely adopted metrics: endpoint-error (EPE), 1-pixel outlier rate (1px), percentage of flow outliers (Fl), and weighted area under the curve (WAUC). Definitions can be found in [41, 31, 9, 34].
5.1 Architecture Details
Input Encoder
We use a frozen ViT-S-based DepthAnythingV2 [60] and a ResNet18 [11] in the input encoder. We only use the feature map at the original image resolution from the DepthAnythingV2 head. This DAv2 feature is concatenated with its corresponding image and then fed into ResNet18 for refinement. We upsample the output of ResNet18 to obtain the final image features at 1/2 image resolution using 3 deconv layers.
Recurrent Update Module
We use a modified DPT-Small [39] as the recurrent update module. We concatenate the image features and use a $1\times 1$ conv to obtain the initial hidden state. Since the image features are already 2×2$\times$2 × downsampled, we change the patch size to 8. We set the resolution of the positional embedding to $224\times 224$, and interpolate it for other resolutions. We use $T=5$ iterations in training and inference.
5.2 Training Details
[…]
FlowNet2 [16] 6.710 1.040 2.823 90.907
SpyNet [40] 29.963 4.162 12.866 67.150
PWC-Net [46] 82.27 2.288 4.889 45.670
RAFT [50] 6.790 1.476 3.198 90.920
GMA [20] 7.074 0.914 3.079 90.722
FlowFormer [13] 6.510 0.723 2.384 91.679
GMFlow [56] 10.355 0.945 2.952 82.337
RPKNet [35] 4.809 0.657 1.756 92.638
CroCoFlow [54] 4.565 0.498 1.508 93.660
Win-Win [24] 5.371 0.475 1.621 92.270
SEA-RAFT(M) [52] 3.686 0.363 1.347 94.534
DPFlow [34] 3.442 0.340 1.311 94.980
WAFT-540p (Ours)
3.418 0.340 1.280 94.663
WAFT-1080p (Ours)
3.347
0.337
1.222
95.189
Table 3: WAFT ranks 1st on Spring [31]. Our 540p model also outperforms existing methods on 1px, EPE, and Fl. denotes the submissions from the Spring team.
Benchmark Submissions
Following SEA-RAFT [52], we first pre-train our model on TartanAir [51] for 300k steps, with a batch size of 32 and learning rate $4\times 10^{-4}$. We fine-tune our model on FlyingChairs [62] with the same hyperparameters for 50k steps, and then fine-tune it on FlyingThings [30] with the same hyperparameters for 200k steps. For all submissions, we keep the batch size as 32 by default and reduce the learning rate to $10^{-4}$. For KITTI [9] submission, we fine-tune our model on KITTI(train) [9] for 5k steps. For Sintel [3] submission, we follow previous work to fine-tune our model on the mixture of FlyingThings [30], HD1K [23], KITTI(train) [9], and Sintel(train) [3] for 200k steps. For Spring [31] submissions, we fine-tune two models on Spring(train) [31] for 200k steps: a 1080p model with a batch size of 8 and a 540p model with a batch size of 32.
Zero-Shot Evaluation
We first train our model on FlyingChairs [8] for 50k steps, and then fine-tune it on FlyingThings [30] for 50k steps. The batch size is set to 32, and the learning rate is set to $10^{-4}$.
5.3 Benchmark Results
Method Sintel (train) KITTI (train)
[…]
PWC-Net [46] 2.55 3.93 10.4 33.7
RAFT [50] 1.43 2.71 5.04 17.4
GMA [20] 1.30 2.74 4.69 17.1
SKFlow [48] 1.22 2.46 4.27 15.5
DIP [66] 1.30 2.82 4.29 13.7
EMD-L [5] 0.88 2.55 4.12 13.5
CRAFT [45] 1.27 2.79 4.88 17.5
RPKNet [35] 1.12 2.45 - 13.0
GMFlowNet [63] 1.14 2.71 4.24 15.4
FlowFormer [13] 1.01 2.40 4.09 14.7
Flowformer++ [44] 0.90 2.30 3.93 14.2
CCMR+ [17] 0.98 2.36 - 12.9
MatchFlow(G) [6] 1.03 2.45 4.08 15.6
SEA-RAFT(L) [52] 1.19 4.11 3.62 12.9
AnyFlow [21] 1.10 2.52 3.76 12.4
SAMFlow [67] 0.87
2.11
3.44 12.3
FlowDiffuser [26]
0.86
2.19 3.61 11.8
DPFlow [34] 1.02 2.26 3.37 11.1
FlowNet2 [16] 2.02 3.14 10.1 30.4
WAFT (Ours)
1.00 2.15
3.10
10.3
Table 4: WAFT achieves the best cross-dataset generalization on KITTI(train), reducing the error by 8%.
Sintel & KITTI
Results are shown in Table 2. On KITTI [9], WAFT achieves the best Fl on non-occluded pixels and the second best on all pixels. WAFT is competitive on Sintel (Final) [3]. On Sintel (Clean), it achieves performance close to state-of-the-art methods. WAFT is up to 4.1×4.1$\times$4.1 × faster than comparable methods, demonstrating its high efficiency.
Spring
Results are shown in Table 3. Benefiting from warping, WAFT can be trained at full resolution (1080p) and achieves the best performance on all 4 metrics. We also train a 540p model following the downsample-upsample protocol of SEA-RAFT [52]; this 540p model outperforms all existing methods on 3 of the 4 metrics: 1px-outlier rate, endpoint-error, and Fl.
Comparison with Existing Warping-based Methods
It appears that warping as an network operation has been largely abandoned by works in the last 8 years and the last time warping-based methods achieved top positions on the leaderboards were around 2017 [40, 16]. WAFT is significant in that it has revisited and revived an idea that has fallen out of favor. Compared to methods that do use warping [40, 16], WAFT reduces endpoint-error (EPE) by at least 59% on Sintel and 67% on Spring [31], while also reducing Fl by at least 67% on KITTI [9] and 56% on Spring [31]. Besides, WAFT reduces 1px-outlier rate by 50% on Spring [31].
5.4 Zero-Shot Evaluation
Following previous work [50, 13, 46], we train our model on FlythingChairs [8] and FlyingThings [30]. Then we evaluate the performance on the training split of Sintel [3] and KITTI [9].
Analysis
Results are shown in Table 4. WAFT achieves strong cross-dataset generalization. On KITTI(train) [9], WAFT outperforms other methods by a large margin: It improves the endpoint-error (EPE) from 3.37 to 3.10 and Fl from 11.1 to 10.3. On Sintel(train) [3], WAFT achieves performance close to state-of-the-art methods. Compared to previous warping-based method [16], WAFT improves the performance by at least 31%.
5.5 Ablation Study
Experiment Input Enc. Rec. Upd. #Steps Index Reso. Sintel(train) Spring(sub-val) #MACs
[…]
WAFT DAv2-S+Res18 DPT-S 5 1/2 1.18 2.33 0.27 1.43 858G
Different input enc. Res18 DPT-S 5 1/2 1.27 2.81 0.27 1.59 600G
DAv2-S 1.55 2.64 0.37 2.70 670G
DAv2 w/o pre-train DAv2-S+Res18 DPT-S 5 1/2 1.42 2.74 0.28 1.77 858G
Different rec. upd. DAv2-S+Res18 Res18 5 1/2 7.23 6.84 0.45 2.93 1098G
ConvGRU 2.79 4.80 0.39 2.71 800G
1/8 reso. + warp DAv2-S+Res18 DPT-S 5 1/8 1.15 2.31 0.32 1.82 859G
1/8 reso. + corr. DAv2-S+Res18 DPT-S 5 1/8 1.10 2.45 0.33 1.74 883G
Direct variants DAv2-B+Res18 DPT-S 1 1/2 2.36 3.43 0.59 10.5 1009G
DPT-B 2.37 3.38 0.61 11.1 1277G
Image-space warp DAv2-S+Res18 DPT-S 5 1/2 1.28 2.50 0.27 1.37 1902G
Refine w/o warp DAv2-S+Res18 DPT-S 5 1/2 2.04 3.42 0.58 9.44 858G
w/ Context DAv2-S+Res18 DPT-S 5 1/2 1.22 2.32 0.29 1.70 1005G
Table 5: We report the zero-shot ablation results on Sintel(train) [3] and Spring(sub-val) [31, 52]. The effect of changes can be identified through comparisons with the first row. See Section 5 for details.
We conduct zero-shot ablations in Table 5 on the training split of Sintel [3] and the sub-val split [52] of Spring [31]. In all experiments, the models are trained on FlythingThings [30] for 50k steps, with a batch size of 32 and learning rate $10^{-4}$. The average end-point-error and 1px-outlier rate are reported.
Different Input Encoder
Both pre-trained DepthAnythingV2 [60] feature and ResNet [11] refinement are important to the model performance.
Different Recurrent Update Module
The vision transformer design is crucial to iterative warping. We observe a significant performance drop when replacing the DPT-based recurrent update module with CNNs, highlighting the importance of modeling long-range dependence. This finding may help explain why early deep learning approaches [16, 40] that implemented warping using CNNs underperformed compared to cost-volume-based methods [46, 50, 13].
High-Resolution Indexing
High-resolution indexing into feature maps using current flow estimates remarkably improves performance. We implement variants that index at 1/8 resolution by changing the patch size of DPT to $2\times 2$, and find that high-resolution indexing significantly improves 1px-outlier rate on Spring(sub-val) [31].
We also design a cost-volume-based variant following the common setup [52, 50, 13] with look-up radius $4$ at 1/8 resolution, and find that it performs similarly to the warping counterpart (shown in Table 5) but costs $\times2$ training memory (21.2 GiB vs. 9.5 GiB).
Direct vs. Iterative
Iterative updates achieve better performance than direct regression within our meta-architecture. We implement direct regression by setting the number of iterations $T=1$. For fair comparison, we scale up the networks to match the computational cost of 5-iteration WAFT. Our results show that WAFT significantly outperforms these direct regression variants, indicating the effectiveness and high efficiency of the iterative paradigm. This finding aligns with the observation that existing direct methods either underperform [56, 53] or require substantially more computational cost [43, 53, 54] compared to iterative approaches [50, 52, 13, 34].
Warping Features vs. Pixels
Feature-space warping is more effective than image-space warping, which is commonly used in classic methods [29, 2, 1] and early deep learning methods [16, 40]. Feature-space warping does not need to re-extract features for the warped image in each iteration, significantly saving computational cost while achieving slightly better accuracy.
Effectiveness of Warping
It is possible to perform iterative updates without warping the features using the current flow estimates. We can simply use the original feature maps as input to the update module. Compared to this baseline, warping has significantly lower error with a negligible cost. This observation aligns with the conclusions of previous work [2], which theoretically justifies the combination of warping and recurrent updates by framing it as a fixed-point iteration algorithm.
Context Encoder
Prior work [46, 50, 52, 13] has often used a context encoder that provides an extra input to the update module. Our ablation show that the context encoder is not necessary. The context encoder introduces additional computation overhead, but does not significantly affect performance. Previous work [52] also points out that the context encoder can be regarded as a direct flow regressor, which functions similarly to the first iteration of WAFT.
6 Conclusion
We have introduced Warping-Alone Field Transforms (WAFT), a simple and effective method for optical flow. WAFT is similar to RAFT but replaces cost volume with high-resolution warping, achieving better accuracy and lower memory cost. This design challenges the conventional wisdom that constructing cost volumes is necessary for strong performance.
7 Acknowledgments
This work was partially supported by the National Science Foundation.
Appendix A Appendix
Sequence WAFT (Ours) DPFlow [34] Flowformer++ [44] FlowDiffuser [26] DDVM [43]
Perturbed Market 3 0.953 (0.416) 0.892 (0.423) 0.958 (0.511) 0.897 (0.490) 0.787 (0.372)
Perturbed Shaman 1 0.208 (0.184) 0.213 (0.201) 0.251 (0.236) 0.267 (0.255) 0.219 (0.196)
Ambush 1 19.40 (3.146) 8.366 (2.970) 6.610 (2.605) 7.224 (2.583) 29.33 (14.07)
Ambush 3 3.669 (2.822) 3.019 (1.729) 2.939 (1.816) 3.148 (1.828) 2.855 (3.016)
Bamboo 3 0.466 (0.446) 0.486 (0.438) 0.546 (0.513) 0.594 (0.508) 0.415 (0.380)
Cave 3 2.260 (1.574) 2.341 (1.631) 2.344 (1.477) 2.464 (1.433) 2.042 (1.658)
Market 1 0.851 (0.434) 0.890 (0.491) 1.073 (0.550) 1.238 (0.517) 0.719 (0.467)
Market 4 6.652 (3.888) 7.939 (3.834) 8.086 (4.450) 8.024 (4.027) 5.517 (3.971)
Mountain 2 0.175 (0.092) 0.177 (0.078) 0.288 (0.118) 0.409 (0.101) 0.176 (0.095)
Temple 1 0.517 (0.332) 0.511 (0.302) 0.657 (0.359) 0.789 (0.340) 0.452 (0.284)
Tiger 0.503 (0.395) 0.571 (0.411) 0.595 (0.430) 0.636 (0.391) 0.413 (0.344)
Wall 2.038 (0.850) 2.011 (1.231) 1.727 (0.793) 1.692 (0.734) 1.639 (3.210)
All 2.337 (1.094) 1.975 (1.046) 1.943 (1.073) 2.026 (1.016) 2.475 (1.754)
Table 6: We report the endpoint-error (EPE) on all sequences of Sintel [3], shown in the format “final-epe (clean-epe)”. We highlight the best result on each sequence.
Sintel Results
We show the sequence results of several representative methods on Sintel [3] in Table 6.